What is the volume in liters occupied by 6.79 g of Freon-12 gas, CCl2F2, at 1.21 atm and 35°C?
1
Expert's answer
2020-08-03T07:41:45-0400

pv = nRT; n = m/M

p = 1.21 atm, m = 6.79 g, M = 2*35.5+2*19+12 = 121 g/mol, R = 8.31, T = 308 K.

v = (m * R * T) / (M * p) = 1.1715 liters.